If alpha and beta are  zeros of the quardric polynomial  6x2+x-2 find the value of alpha/beta + beta/alpha

Grade:10

1 Answers

Arun
6 years ago
let roots are a & b.
then we have to find the value of a/b + b/a
now,
= (a^2+b^2)/ab
= [(a+b)^2-2ab]/ab
but we know that sum of roots = a+b = -1/6 and product of roots =ab= 1/3
so now,
= (1/36−2/3)/(1/3)
−23/12